A particle is moving eastward with a velocity of 5 m/s into 10 second velocity changes to 5 m/s NorthWord average acceleration in metre per second Square in this time is

A particle  is moving eastward with a velocity of 5 m/s into 10 second velocity changes to 5 m/s NorthWord average acceleration in metre per second Square in this time is 

Grade:11

1 Answers

Issacc Jayan
6 years ago
Since velocity is a vector quantity, by triangle law of vector addition, resultant vector equals 5×1.414(root 2)= 7.07m/s. Acceleration =7.07/10 =.707m/s^2